Steering Manufacturing produces snow shovels. The selling price per snow shovel is $29.00. There is no beginning inventory.

Costs involved in production are:
Direct material $5.00
Direct labor 4.00
Variable manufacturing overhead 4.00
Total variable manufacturing costs per unit $13.00
Fixed manufacturing overhead per year $181,170


In addition, the company has fixed selling and administrative costs of $172,400 per year.

During the year, Steering produces 54,900 snow shovels and sells 49,560 snow shovels.

What is the value of ending inventory using full costing?

Value of ending inventory $_______________
Answer #1

Calculate unit product cost under full costing

Direct material 5
Direct labor 4
Variable manufacturing overhead 4
Fixed manufacturing cost (181170/54900) 3.30
Unit product cost 16.30

Ending inventory unit = 54900-49560 = 5340

value of ending inventory = 5340*16.30 = $87042